ganeshie8 (ganeshie8):

[Weekend Challenge - Solved by @adxpoi and @oldrin.bataku ] Suppose \(P(x)\) is a polynomial of degree \(10\) such that \(P(k) = 2^k\) for all \(k=0,1,2,\ldots,10\). Find the value of \(P(11)\)

OpenStudy (freckles):

\[P(x)=a_{10}x^{10}+a_9x^9+a_8x^8+a_7x^7+a_6x^6+a_5x^5+a_4x^4+a_3x^3 + \\ +a_2x^2+a_1x^+a_0 \\ P(0)=1 \\ a_0=1 \\ \] well I know there is a really long way... pluggin in all those critters to find the constant coefficients of the polynomial P

OpenStudy (anonymous):

4095?

ganeshie8 (ganeshie8):

11 equations and 11 unknowns is solvable for sure :)

ganeshie8 (ganeshie8):

@ospreytriple nope, but it seems you got it! the answer is \(2047\)

ganeshie8 (ganeshie8):

please post your method, im pretty sure it is just a typo somewhere..

OpenStudy (anonymous):

Darn. Added an extra factor of 2 somewhere. :)

OpenStudy (anonymous):

Not sure if I understood the question but the condition that \(P(k)=2^k\) got me thinking that the answer was simply \(P(11)=2^{11}+2^{10}+...+2^1+2^0

OpenStudy (triciaal):

why not 2048?

OpenStudy (triciaal):

never mind

ganeshie8 (ganeshie8):

Because we only know \(11\) values of function \(P(x)\), namely, \(P(0), P(1), P(2),\ldots,P(10)\) we don't know the value of\(P(x)\) for other values of \(x\)

OpenStudy (anonymous):

Consider the polynomial \[Q(x) = \sum_{j=0}^{10} (-1)^{10-j}\frac{2^j}{j!(10-j)!}l_j(x)\] where \(l_j(x) = \frac{x(x-1)...(x-10)}{(x-j)}\) [\(Q(x)\) is basically the Lagrange Interpolation polynomial with data points \((0,1),(1,2),(2,2^2),...,(10,2^{10})\)] Note that \(l_j(j) = (-1)^{10-j}j!(10-j)!\) Therefore, \(Q(j) = 2^j\) for \(j = 0\) to \(10\) \(Q(x)\) is a polynomial of degree at most \(10\) so, \(Q(x)-P(x)\) is also a polynomail of degree at most \(10\) Also, \(Q(x)-P(x) = 0\) for \(11\) values of \(x\). Therefore, \(Q(x) -P(x)\) ia identically zero. That is, \(Q(x) = P(x)\) Therefore, \(P(11)\) \( = Q(11)\) \( = \sum_{j=0}^{10} (-1)^{10-j}\frac{2^j}{j!(10-j)!}l_j(11)\) \( = \sum_{j=0}^{10} (-1)^{10-j}\frac{2^j}{j!(10-j)!}\frac{11!}{(11-j)}\) \( = \sum_{j=0}^{10} (-1)^j2^j\left(\begin{matrix}11 \\ j\end{matrix}\right)\) \( = 2^{11} - (2-1)^{11}\) \( = 2^{11} - 1\)

OpenStudy (anonymous):

@adxpoi did it correctly, although I used Newton polynomials

OpenStudy (anonymous):

consider the finite differences: 1 1 2 1 2 4 2 4 8 4 8 16 8 16 32 which is a consequence of the fact that \(2^k=2\cdot2^{k-1}=2^{k-1}+2^{k-1}\). now consider $$ P(n)=\sum_{k=0}^{10}\binom{n}{k}\Delta^k[P](0)\\P(11)=\sum_{k=0}^{10}\binom{11}k\cdot 1=\sum_{k=0}^{10}\binom{11}k $$ now using the binomial theorem, we have that: $$2^{11}=\left(1+1\right)^{11}=\sum_{k=0}^{11}\binom{11}k\\2^{11}-1=\sum_{k=0}^{10}\binom{11}k$$ giving us \(P(11)=2^{11}-1=2047\)

OpenStudy (anonymous):

alternatively, if you understand that \(2^n\) has identical finite differences, i.e. \(\Delta^k[2^n]=2^n\), then we should expect that for our \(10\)th degree polynomial once we reach our \(10\)th difference we'll hit all constants, which must be \(1,1,\dots\). however, for \(2^n\) it would have been changing again, so \(1,2,\dots\).

OpenStudy (anonymous):

so for the \(11\)th term, in which we'll make use of our eleventh difference at \(0\), we expect \(\Delta^{11}[P]=0\), while for \(2^n\) we'd have \(\Delta^{11}[2^n]=1\). hence our result will be \(1\) less than\(2^11\)

ganeshie8 (ganeshie8):

Excellent! both methods look great! In summary, the polynomial that satisfies given data points is \[\large P(x) = \sum\limits_{i=0}^{10} \binom{x}{i}\] plugging in \(x=11\) gives \(P(11) = 2^{11}-1\)

OpenStudy (anonymous):

I tried the brute force method. With\[P(x) = a_{10}x^{10} + a_9x^9+a_8x^8+a_7x^7+a_6x^6+a_5x^5+a_4x^4+a_3x^3+a_2x^2+a_1x+a_0\]and knowing that\[P(0)=1\]\[P(1)=2\]\[P(2)=4\]\[P(3)=8\]\[P(4)=16\]\[P(5)=32\]\[P(6)=64\]\[P(7)=128\]\[P(8)=256\]\[P(9)=512\]\[P(10)=1024\]I created an augmented matrix and after a lot of Gauss-Jordan elimination determined that\[a_{10}=2.75576\times10^{-7}\]\[a_9=-9.64519\times10^{-6}\]\[a_8=1.65347\times10^{-4}\]\[a_7=-1.59560\times10^{-3}\]\[a_6=1.01449\times10^{-2}\]\[a_5=-3.87451\times10^{-2}\]\[a_4=1.12450\times10^{-1}\]\[a_3=-1.05290\times10^{-1}\]\[a_2=3.77246\times10^{-1}\]\[a_1=6.45634\times10^{-1}\]\[a_0=1\]I kept about double these number of decimal places and confirmed the results using Excel. After confirming the value of the coefficients, I was able to calculate\[P(11)=2047\]I'm tired now :)